# Algebraic Topology Principles behind Topological Quantum Error Correction

### v1

@@ Abstract
Quantum error correction (QEC) is crucial for numerous quantum applications, including fault-tolerant quantum computation, which is of great scientific and industrial interest. Among various QEC paradigms, topological quantum error correction (TQEC) has attained the most experimental successes by far. In this paper, we build upon existing knowledge of TQEC by developing a generalized theoretical framework of TQEC. We begin by formally defining TQEC codes and exploring the algebraic topological principles underlying these quantum codes, including deriving the conditions for any topological manifold to serve as a quantum memory. We show that TQEC for qubits works for both orientable and non-orientable manifolds. Moreover, we extend the construction of TQEC to higher-dimensional manifolds and provide examples for higher-dimensional TQEC codes. Finally, we apply these principles to construct new codes on 2-dimensional manifolds that have received limited attention in prior literature. As a case study, we simulate the performance of TQEC codes on the Klein bottle $K$ and evaluate their efficacy for quantum error correction. This work contributes to the advancement of TQEC by proposing a broader class of codes and demonstrating their theoretical and practical potential. By addressing previously unexplored topological structures, our findings represent a step forward in achieving fault-tolerant quantum computation and communication.
@@ I Introduction
The concept of topological quantum error correction (TQEC) was initially introduced by Kitaev [21], wherein two qubits are encoded on the two-dimensional topological manifold of a torus ($T^{2}=S^{1}\times S^{1}$). This pioneering framework laid the foundation for subsequent advancements in TQEC codes. In pursuit of practical implementations, several later studies have focused on models of TQEC that are suitable to experimental realization. Notably, surface codes defined on planar geometries with boundaries, commonly referred to as planar codes, have been extensively investigated. These studies have shown that TQEC codes can be effectively realized on planar surfaces with non-contractible loops, thereby extending the applicability of the original toric code framework [21, 23, 24]. In parallel, Bombin and Martin-Delgado introduced a comprehensive theory of homological quantum error correction [24]. Their work rigorously established that non-orientable surfaces are unsuitable for TQEC codes defined on qudits, where a qudit is a quantum system with a finite-dimensional Hilbert space of dimension $D>2$.
@@
In this paper, we shall only focus on the TQEC codes for qubits. In Section II, we give an overall review on algebraic topology and the topological properties of manifolds that are crucial for later discussions. Then in Section III, we formulate the theory of TQEC codes with the intersections between homology and cohomology loops, propose extending the TQEC paradigm beyond traditional topological manifolds, and use different manifolds that are seldom discussed, including the non-orientable maniolds like the Klein bottle.
@@
Furthermore, in Section IV, we explore the extension of TQEC into higher-dimensional manifolds that satisfy specific conditions. We believe that an $n$-dimensional manifolds $M$ can encode qubits on the $i$-th dimension if the $i$-th homology group $H_{i}(M;\mathbb{Z}_{2})$ is non-trivial, and we further propose an implementation for a 3-dimensional code. These higher-dimensional structures encompass a broader class of TQEC codes, which hold promise for future applications and are of substantial theoretical interest in advancing the understanding of TQEC principles.
Thus, we address the following theoretical aspects in the paper:
  1. 1.
    The topological properties for a manifold to be used for TQEC.
  2. 2.
    The generalization of TQEC to higher-dimensional manifolds, and their implementation.
  3. 3.
    The practicality and advantages of using 2-manifolds other than the torus for TQEC.
@@
Finally, in Section V, we demonstrate the simulation result of the TQEC codes using the Klein Bottle, which was seldom discussed in the literature. The result gives great potential for developing new TQEC codes for future applications.
@@ II Topological Properties of Manifolds
Before discussing the homology, we first review the standard definition of the boundary and coboundary maps. Denote the set of chains complex on a manifold $M$ over a commutative ring $R$ as $C_{n}(M;R)=R\langle\sigma:\Delta^{n}\rightarrow M\rangle$, where $\Delta^{n}$ represents the $n$-simplex on $M$ [26, 27], the formal definition of simplexes is in Appendix A.2. Then, we can define the boundary map, $\partial_{n}:C_{n}\rightarrow C_{n-1}$, where $\partial_{n}$ takes an $n$-simplex from $C_{n}(M)$ and return a $(n-1)$-simplex which is the boundary of the $n$-simplex [26, 27]. Similarly, we can define the co-boundary map [24], $\delta^{i}:C^{i-1}\rightarrow C^{i}$, which is the dual homomorphism of $\partial_{i}$. This means, for some cell complex $X_{M}$, and some co-chain $c^{n}\in C^{n}(X_{M};R)$, and some chain $c_{n}\in C_{n}(X_{M};R)$ [24]
@@
[…]
@@
[…]
@@
$H^{n}(X)=Z^{n}(X;R)/B^{n}(X;R)$
@@
This corollary implies that for any manifolds $M$, and any cell complex $X_{M}$ on $M$, the singular chain complex $C_{i}(M;R)$ and the cell chain complex $C_{i}^{cell}(X_{M};R)$ are isomorphic as chain complexes. And then, because of this chain complex isomorphism, the singular homology $H(M;R)$ is isomorphic to the simplicial homology $H^{cell}(X_{M};R)$. Therefore, by choosing different cell complexes on the manifold $M$, the homology is invariant. As we will later prove in section III.3, the TQEC ability of a surface is independent of the cell complex we choose to embed on the manifold.
@@
Theorem II.1.
(Poincaré duality) For any commutative ring $R$, if $M$ is a closed $R$-orientable $n$-manifold with fundamental class $[M]\in H_{n}(M;R)$, then the map $D:H^{k}(M;R)\rightarrow H_{n-k}(M;R)$ defined by $D(\alpha)=[M]\cap\alpha$ is an isomorphism for all $k$.
The above definition uses the cap product $\cap$, formally defined in Appendix A.3. Notice that the above definition involves homology groups with coefficients from a coefficient ring $R$, where $R$ is usually taken as $\mathbb{Z}$. Then the theorem II.1 states that for any compact oriented $n$-manifolds, there is an isomorphism between $H_{k}(M)$ and $H^{n-k}(M)$. However, since we are interested in using a topological manifold as TQEC codes for qubits, for the majority of this paper, the coefficient ring $R$ is considered to be $\mathbb{Z}_{2}$ as mentioned above. The condition of the surface becomes $\mathbb{Z}_{2}$-orientable, where, by definition, all compact manifolds, including the non-orientable manifolds, are $\mathbb{Z}_{2}$-orientable [29, 30].
@@
The relationship between the homology and cohomology groups can be regarded as the unique pairing between the $[\gamma]\in C_{i}(M)$ and $[\beta]\in C^{i}(M)$, which can be interpreted as the number of intersections. For a manifold, given a $i$-chain $\gamma\in C_{i}(M,\mathbb{Z}_{2}),\partial\gamma=0$ and $[\gamma]\neq 0$ and a class of functionals on loops $\beta\in C^{i}(M,\mathbb{Z}_{2}),\delta\beta=0$ and $[\beta]\neq 0$, we can “evaluate” the functionals on the loop gives a number in $\mathbb{Z}_{2}=\{0,1\}$. This can be interpreted as the intersection points between $[\gamma]$ and $[\beta]$. Poincaré duality in equation II.1 has non-trivial implications for the intersections between homology and cohomology groups [31, 32], that for any closed and oriented n-manifold $M$. The Poincaré duality can be reformulated as,
@@ III.1 Definition of TQEC Codes on 2-Manifolds
With the properties of manifolds established, we proceed to define topological quantum error correction (TQEC) codes on cell complexes embedded in 2-manifolds, which will subsequently be extended to higher-dimensional manifolds. While the formal definition of a cell complex is provided in Appendix A.2, for simplicity, it can be conceptualized as decomposing a two-dimensional manifold into constituent parts, formed by vertices, edges, and faces, and organizing these components into corresponding sets. In this framework, we focus on quantum error correction codes for qubits [21].
@@
  • Vertex Operator: $A_{i}=\sigma_{Z}\left(\delta v_{i}\right)=\sigma_{Z}\left(\partial\bar{f_{i}}\right)$
  • Plaquette Operator: $B_{j}=\sigma_{X}\left(\partial f_{j}\right)$
In simple terms, the vertex operators $A_{i}$ are defined as the stabilizer formed by the tensor product of $\sigma_{Z}$ of qubits on edges with boundary $v$; and the plaquette operators $B_{j}$ are defined as the stabilizer formed by the tensor product of $\sigma_{X}$ at the boundary of each face $f$. It is easily checkable that $\Pi_{v}A_{v}=\Pi_{p}B_{p}=\mathbb{I}$, and all the stabilizers commute, where $\mathbb{I}$ is the identity. And, we define the code space of our TQEC to be the ground state of the following Hamiltonian, for some constant $J$:
@@
A logical $\sigma_{X}$ is defined to operate on any cycles $\gamma\in C_{1}(M;\mathbb{Z}_{2})$, such that $\gamma\subset E$, where $\partial[\gamma]=0$ and $[\gamma]\neq 0$. Similarly, a logical $\sigma_{Z}$ is defined to operate on any co-cycles, $\beta\in C^{1}(M;\mathbb{Z}_{2})$, such that $\beta\subset E$, where $\delta[\beta]=0$ and $[\beta]\neq 0$.
Thus, the above definition of logical $\sigma_{X}$ and $\sigma_{Z}$ satisfies the properties of Pauli operations. Firstly, by the definition of $\mathbb{Z}_{2}$ coefficient, $[\gamma]+[\gamma]=0$ and $[\beta]+[\beta]=0,\forall\gamma,\beta$, meaning that operate logical $\sigma_{X}^{2}=\sigma_{Z}^{2}=\mathbb{I}$. Also, using the pairing property as presented in equation 7, $[\beta]([\gamma])=1$, meaning that $[\beta]$ and $[\gamma]$ intersect at an odd number of sites, thus, the logical $\sigma_{X}$ and $\sigma_{Z}$ anti-commutes, the idea is illustrated in Figure 1.

### v2

@@ Abstract
Quantum error correction (QEC) is crucial for realizing scalable quantum technologies, and topological quantum error correction (TQEC) has emerged as the most experimentally advanced paradigm of QEC. Existing homological and topological code constructions, however, are largely confined to orientable two-manifolds with simple boundary conditions, and higher-dimensional or non-manifold generalizations are typically treated on a case-by-case basis rather than within a single unifying framework. In this work, we develop a unified algebraic-topological framework for TQEC based on homology, cohomology, and intersection theory, which characterizes exactly when an arbitrary-dimensional manifold (with or without boundary) can serve as a quantum memory, thereby extending the standard 2D homological-code picture to arbitrary dimension and to manifolds with boundary via Poincaré-Lefschetz duality. Building on this classification, we introduce concrete code families that exploit nontrivial topology beyond the planar and toric settings. These include “3-torus code” and higher-dimensional “volume codes” on compact manifolds with mixed $X$- and $Z$-type boundaries. We further give a topological construction of qudit TQEC codes on general two-dimensional cell complexes using group presentation complexes, which unifies and extends several known quantum LDPC and homological-product-like constructions within a single geometric language. Finally, we combine the theoretical framework with numerical simulations to demonstrate that changing only the global topology - for example, from a torus to a Klein bottle - can yield improved logical performance at fixed entanglement resources. Taken together, our results provide a systematic set of topological design principles for constructing and analyzing TQEC codes across dimensions, boundaries, and local dimensions, and they open new avenues for topology-aware fault-tolerant quantum architectures.
@@ I Introduction
The concept of topological quantum error correction (TQEC) was initially introduced by Kitaev [21], wherein two qubits are encoded on the two-dimensional topological manifold of a torus ($T^{2}=S^{1}\times S^{1}$). This pioneering framework laid the foundation for subsequent advancements in TQEC codes. In pursuit of practical implementations, several later studies have focused on models of TQEC that are suitable for experimental realization. Notably, surface codes defined on planar geometries with boundaries, commonly referred to as planar codes, have been extensively investigated. These studies have shown that TQEC codes can be effectively realized on planar surfaces with non-contractible loops, thereby extending the applicability of the original toric code framework [21, 23, 24]. In parallel, Bombin and Martin-Delgado introduced a comprehensive theory of homological quantum error correction [24]. Their work rigorously established that non-orientable surfaces are unsuitable for TQEC codes defined on qudits, where a qudit is a quantum system with a finite-dimensional Hilbert space of dimension $D>2$.
@@
In this paper, we first focus on the theory behind qubit codes. In Section II, we give an overall review of algebraic topology and the topological properties of manifolds that are crucial for our discussions. Then in Section III, we formulate the theory of TQEC codes with the intersections between homology and cohomology loops, propose extending the TQEC paradigm beyond traditional topological manifolds, and use different manifolds that are seldom discussed, including the non-orientable manifolds like the Klein bottle.
@@
Furthermore, in Section IV, we explore the extension of TQEC into higher-dimensional manifolds that satisfy specific conditions. We believe that an $n$-dimensional manifold $M$ can encode qubits on the $i$-th dimension if the $i$-th homology group $H_{i}(M;\mathbb{Z}_{2})$ is non-trivial, and we further propose an implementation for a 3-dimensional code. These higher-dimensional structures encompass a broader class of TQEC codes, which hold promise for future applications and are of substantial theoretical interest in advancing the understanding of TQEC principles.
Subsequently, in Section V, we extend the theory to include TQEC codes on manifolds with boundary, which involves the ideas of relative cycles and relative homology reviewed in Section II, which then allows us to construct TQEC codes on manifolds with boundary in any arbitrary dimensions, analogous to the “surface code” in 2-dimensions.
Then, we shift our focus from qubit codes to qudit codes in Section VI, where we demonstrate a bottom-up construction of the TQEC code on a 2-dimensional cell-complex for a general $d$-dimensional qudit, which is of great theoretical interest. We further showed that some quantum Low Parity Density Check (qLDPC) codes are special cases of our code construction.
@@
Finally, in Section VII, we demonstrate the simulation results of the TQEC codes using the Klein Bottle and $\mathbb{R}P^{2}$, which were seldom discussed in the literature. The result gives great potential for developing new TQEC codes for future applications.
As we are aware, our work is among the first few works that generalize the theory of TQEC to arbitrary-dimensional manifolds with and without boundaries, which represents a great step forward in understanding the theory behind TQEC and achieving fault-tolerant quantum computation and communication.
@@ II Topological Properties of Manifolds
II.1 Chain Complexes, Relative Chain Complexes and Cycles
Before discussing the homology, we first review the standard definition of the boundary and coboundary maps. Denote the set of chains complex on a manifold $M$ over a commutative ring $R$ as $C_{n}(M;R)=R\langle\sigma:\Delta^{n}\rightarrow M\rangle$, where $\Delta^{n}$ represents the $n$-simplex on $M$ [26, 27], the formal definition of simplexes is in Appendix B.2. Then, we can define the boundary map, $\partial_{n}:C_{n}\rightarrow C_{n-1}$, where $\partial_{n}$ takes an $n$-simplex from $C_{n}(M)$ and return a (n1)(n-1)-simplex which is the boundary of the $n$-simplex [26, 27]. Similarly, we can define the co-boundary map [24], $\delta^{i}:C^{i-1}\rightarrow C^{i}$, which is the dual homomorphism of $\partial_{i}$. This means, for some cell complex $X_{M}$, and some co-chain $c^{n}\in C^{n}(X_{M};R)$, and some chain $c_{n}\in C_{n}(X_{M};R)$ [24]
@@ II.1 Chain Complexes, Relative Chain Complexes and Cycles
Given a topological space $X$, one typically takes $C_{n}(X)$ to be the free abelian group generated by singular $n$-simplices in $X$. For a subspace $A\subseteq X$, the relative chain complex $C_{n}(X,A)$ is defined as the quotient $C_{n}(X)/C_{n}(A)$, whose boundary maps are induced from those of $C_{n}(X)$. The relative chain complex thus measures chains in $X$ modulo those that lie entirely in $A$ [26, 27].
@@
$\begin{split}\text{$n$-cycle in X: }Z_{n}(X;R)&:=\text{Ker}(\partial_{n})\\ \text{$n$-boundary in X: }B_{n}(X;R)&:=\operatorname{Im}(\partial_{n+1})\end{split}$ (2)
In a chain complex $(C_{*},\partial_{*})$, an element $z\in C_{n}$ is called an $n$-cycle if it lies in the kernel of the boundary operator, that is, $\partial_{n}(z)=0$. The group of all $n$-cycles is denoted by $Z_{n}=\ker(\partial_{n})$. Similarly, an element $b\in C_{n-1}$ is called an (n1)(n-1)-boundary if it lies in the image of the boundary map from the next chain group, that is, $b=\partial_{n}(c)$ for some $c\in C_{n}$; the group of all such boundaries is denoted by $B_{n-1}=\operatorname{Im}(\partial_{n})$. Because $\partial_{n-1}\circ\partial_{n}=0$, every boundary is a cycle, hence $B_{n}\subseteq Z_{n}$.
For a pair of topological spaces $(X,A)$ with $A\subseteq X$, one can define the corresponding relative chain complex:
@@
The relative chain cycle of a manifold $X$ relative to a sub-manifold $A\subseteq X$ is defined as
$C_{n}(X,A)=C_{n}(X)/C_{n}(A)$ (3)
with induced boundary maps $\partial_{n}:C_{n}(X,A)\to C_{n-1}(X,A)$. An element $z]\in C_{n}(X,A)$ (the equivalence class of $z\in C_{n}(X)$) is called a relative $n$-cycle if $\partial_{n}([z])=0$ in $C_{n-1}(X,A)$, that is, if $\partial_{n}z\in C_{n-1}(A)$. The group of relative $n$-cycles is therefore given by
$Z_{n}(X,A)=\{z\in C_{n}(X)\mid\partial_{n}z\in C_{n-1}(A)\},$ (4)
and the group of relative $n$-boundaries by
$B_{n}(X,A)=\{\partial_{n+1}c\mid c\in C_{n+1}(X),\ \partial_{n+1}c\in C_{n}(X)\}.$ (5)
Which means, relative $n$-cycles $z\in Z_{n}(X,A)$ are chain in $C_{n}(X)$ with their boundary in $C_{n-1}(A)$. The concept of relative cycles is crucial when we discuss TQEC codes on manifolds with boundary and holes in Section V.
II.2 Homology and Relative Homology
Definition II.2.
@@ II.2 Homology and Relative Homology
For a pair $(X,A)$ consisting of a topological space $X$ and a subspace $A\subseteq X$, the relative chain complex is given by the quotient $C_{*}(X,A)=C_{*}(X)/C_{*}(A)$. The corresponding boundary operators induce well-defined maps on the quotient. The relative homology groups are then defined by
$H_{n}(X,A)=H_{n}(C_{*}(X,A))=\operatorname{Ker}(\partial^{\prime}_{n})/\operatorname{Im}(\partial^{\prime}_{n+1}),$ (7)
which measure how the topology of $X$ differs from that of its subspace $A$, capturing information about the structure of $X$ relative to $A$. The concept of relative homology will become crucial when we discuss TQEC codes on compact manifolds with boundary and holes.
II.3 Cohomology and Relative Cohomology
@@
$\begin{split}\text{$n$-cocycle in X: }Z^{i}(X;R)&:=\text{ker}(\delta^{n})\\ \text{$n$-coboundary in X: }B^{i}(X;R)&:=\text{Im}(\delta^{n-1})\end{split}$ (9)
@@
$H^{n}(X)=Z^{n}(X;R)/B^{n}(X;R)$ (10)
For a pair $(X,A)$ of topological spaces with $A\subseteq X$, the relative cochain complex is defined as the subcomplex $C^{*}(X,A;G)=\operatorname{Ker}(C^{*}(X;G)\to C^{*}(A;G))$, where the map is induced by restriction of cochains. The coboundary maps descend naturally to this subcomplex, giving a well-defined cochain complex whose cohomology groups are the relative cohomology groups.
$H^{n}(X,A;G)=H^{n}(C^{*}(X,A;G))=\operatorname{Ker}(\delta^{n})/\operatorname{Im}(\delta^{n-1}).$ (11)
Intuitively, $H^{n}(X,A;G)$ measures the cohomological information of $X$ that vanishes on $A$, or equivalently, the obstructions to extending cochains on $A$ to cochains on $X$.
@@
This corollary implies that for any manifold $M$, and any cell complex $X_{M}$ on $M$, the singular chain complex $C_{i}(M;R)$ and the cell chain complex $C_{i}^{cell}(X_{M};R)$ are isomorphic as chain complexes. And then, because of this chain complex isomorphism, the singular homology $H(M;R)$ is isomorphic to the simplicial homology $H^{cell}(X_{M};R)$. Therefore, by choosing different cell complexes on the manifold $M$, the homology is invariant. As we will later prove in section III.3, the TQEC ability of a surface is independent of the cell complex we choose to embed on the manifold.
II.4 Poincaré Duality and Poincaré–Lefschetz isomorphisms
@@
Theorem II.1 (Poincaré duality).
For any commutative ring $R$, if $M$ is a closed $R$-orientable $n$-manifold with fundamental class $M]\in H_{n}(M;R)$, then the map $D:H^{k}(M;R)\rightarrow H_{n-k}(M;R)$ defined by $D(\alpha)=[M]\cap\alpha$ is an isomorphism for all $k$.
The above definition uses the cap product $\cap$, formally defined in Appendix B.4. Notice that the above definition involves homology groups with coefficients from a coefficient ring $R$, where $R$ is usually taken as $\mathbb{Z}$. Then the theorem II.1 states that for any compact oriented $n$-manifolds, there is an isomorphism between $H_{k}(M)$ and $H^{n-k}(M)$. However, since we are interested in using a topological manifold as TQEC codes for qubits, for much of this paper, the coefficient ring $R$ is considered to be $\mathbb{Z}_{2}$ as mentioned above. The condition of the surface becomes $\mathbb{Z}_{2}$-orientable, where, by definition, all compact manifolds, including the non-orientable manifolds, are $\mathbb{Z}_{2}$-orientable [29, 30].
@@
The relationship between the homology and cohomology groups can be regarded as the unique pairing between the $\gamma]\in C_{i}(M)$ and $\beta]\in C^{i}(M)$, which can be interpreted as the number of intersections. For a manifold, given a $i$-chain $\gamma\in C_{i}(M,\mathbb{Z}_{2}),\partial\gamma=0$ and $\gamma]\neq 0$ and a class of functionals on loops $\beta\in C^{i}(M,\mathbb{Z}_{2}),\delta\beta=0$ and $\beta]\neq 0$, we can “evaluate” the functionals on the loop gives a number in $\mathbb{Z}_{2}=\{0,1\}$. This can be interpreted as the intersection points between $\gamma]$ and $\beta]$. Poincaré duality in equation II.1 has non-trivial implications for the intersections between homology and cohomology groups [31, 32], for any closed and oriented n-manifold $M$. The Poincaré duality can be reformulated as,
@@
When discussing the manifolds with boundary, the usual Poincaré duality does not apply. Therefore, here we introduce a generalization of Poincaré duality [26, 27],
Theorem II.2 (Poincaré–Lefschetz duality).
Let $M$ be a compact, $R$-orientable $n$ manifold with boundary $\partial M$, and let $R$ be a commutative ring. For each $0\leq k\leq n$, the cap product with the relative fundamental class induces natural isomorphisms
$D_{M}\colon H^{k}(M;R)\xrightarrow{\ \ \frown[M,\partial M]\ \ }H_{n-k}(M,\partial M;R)$ (15)
$D_{M}^{\partial}\colon H^{k}(M,\partial M;R)\xrightarrow{\ \ \frown[M,\partial M]\ \ }H_{n-k}(M;R).$ (16)
If $\partial M=\varnothing$, these reduce to the classical Poincaré duality isomorphisms $H^{k}(M;R)\cong H_{n-k}(M;R)$.
The Poincaré-Lefschetz duality induced the following corollary, which is stated as Theorem 3.43 from Ref. [26]
Corollary II.2.1.
Given a compact $R$ orientable $n$ manifold $M$ with boundary $\partial M$, if $\partial M$ is further decomposed as the union of two compact (n1)(n-1) dimensional manifolds $A$ and $B$ with a common boundary $\partial A=\partial B=A\cap B$. Then the cap product with a fundamental class $M]\in H_{n}(M,\partial M;R)$ gives isomorphisms $D_{M}:H^{k}(M,A;R)\longrightarrow H_{n-k}(M,B;R)$ for all $k$.
It is clear to see that Theorem II.2.1 will reduce to the Poincaré-Lefschetz duality in Theorem II.2 when $A=\emptyset$ or $B=\emptyset$; this then reduces to the usual Poincaré duality when $\partial M=\emptyset$. The duality in Theorem II.2.1 will be useful when we consider TQEC codes on compact manifolds with boundaries.
@@ III.1 Definition of TQEC Codes on 2-Manifolds
With the properties of manifolds established, we proceed to define topological quantum error correction (TQEC) codes on cell complexes embedded in 2-manifolds, which will subsequently be extended to higher-dimensional manifolds. While the formal definition of a cell complex is provided in Appendix B.2, for simplicity, it can be conceptualized as decomposing a two-dimensional manifold into constituent parts, formed by vertices, edges, and faces, and organizing these components into corresponding sets. In this framework, we focus on quantum error correction codes for qubits [21].
@@
  • Vertex Operator:
    $A_{i}=\sigma_{Z}\left(\delta v_{i}\right)=\sigma_{Z}\left(\partial\bar{f_{i}}\right)$ (17)
  • Plaquette Operator:
    $B_{j}=\sigma_{X}\left(\partial f_{j}\right)$ (18)
In simple terms, the vertex operators $A_{i}$ are defined as the stabilizer formed by the tensor product of $\sigma_{Z}$ of qubits on edges with boundary $v$; and the plaquette operators $B_{j}$ are defined as the stabilizer formed by the tensor product of $\sigma_{X}$ at the boundary of each face $f$. It is easily checkable that $\Pi_{v}A_{v}=\Pi_{p}B_{p}=\mathbb{I}$, and all the stabilizers commute, where $\mathbb{I}$ is the identity. Figure 1 shows a well-established example of defining the two types of stabilizers on a 2-dimensional lattice used by the Toric code [21].
And, we define the code space of our TQEC to be the ground state of the following Hamiltonian, for some constant $J$:
@@
A logical $\sigma_{X}$ is defined to operate on any cycles $\gamma\in C_{1}(M;\mathbb{Z}_{2})$, such that $\gamma\subset E$, where $\partial[\gamma]=0$ and $\gamma]\neq 0$. Similarly, a logical $\sigma_{Z}$ is defined to operate on any co-cycles, $\beta\in C^{1}(M;\mathbb{Z}_{2})$, such that $\beta\subset E$, where $\delta[\beta]=0$ and $\beta]\neq 0$.
Thus, the above definition of logical $\sigma_{X}$ and $\sigma_{Z}$ satisfies the properties of Pauli operations. Firstly, by the definition of $\mathbb{Z}_{2}$ coefficient, $\gamma]+[\gamma]=0$ and $\beta]+[\beta]=0,\forall\gamma,\beta$, meaning that operate logical $\sigma_{X}^{2}=\sigma_{Z}^{2}=\mathbb{I}$. Also, using the pairing property as presented in equation 14, $\beta]([\gamma])=1$, meaning that $\beta]$ and $\gamma]$ intersect at an odd number of sites, thus, the logical $\sigma_{X}$ and $\sigma_{Z}$ anti-commutes, the idea is illustrated in Figure 2.